# Recent Advances in Attack and Defense Approaches of Large Language Models

### v2

@@ 3.1.2 Alignment Attacks
Alignment attacks can be broadly categorized into two types. The first category is algorithmic attacks, which exploit vulnerabilities inherent in the alignment algorithms themselves. These attacks aim to undermine the integrity of the alignment process by directly targeting the algorithms’ weaknesses. The second category is data poisoning attacks, which focus on corrupting the training data used in the alignment process. A significant portion of research in this area concentrates on reward hacking (Skalse et al., 2022; Shi et al., 2023), where adversaries manipulate the reward mechanisms to achieve undesired outcomes. By tampering with the data that shapes the model’s behavior, these attacks can lead to misalignment and compromise the system’s intended functionality.
@@
Widely adopted alignment methods include reward-guided Proximal Policy Optimization (PPO) (Schulman et al., 2017) and reward-free Direct Preference Optimization (DPO), with DPO considered a more efficient alternative to PPO. However, Pathmanathan et al. (2024) conduct an empirical study revealing that both methods are vulnerable to backdoor and non-backdoor attacks, with DPO being more susceptible across a range of LLMs compared to PPO. Unlike PPO-based methods—which require at least 4% of the data to be poisoned to trigger harmful behavior—DPO can be compromised with as little as 0.5% of poisoned data. Furthermore, Lee et al. (2024) perform a case study to investigate the underlying mechanisms of the DPO algorithm. They discovered that while DPO does not eliminate the generation of toxic outputs, it instead avoids regions that produce toxicity by learning an "offset" distributed across model layers. Based on these findings, they propose a method to reactivate the toxicity of aligned models. Both studies highlight the vulnerabilities in alignment mechanisms and the inadequacy of current defense strategies against these weaknesses.
@@
Another branch of research identifies the reward model as a new attack surface, where data poisoning is particularly effective and stealthy against current defenses. Shi et al. (2023) demonstrate that backdoor attacks can evade detection, causing the reward model to assign high scores to incorrect sentiment classes when a trigger appears, severely impacting the LLM’s performance on sentiment tasks trained with this poisoned reward model. This threat model is further examined in Baumgärtner et al. (2024), which shows that reward data poisoning can be highly effective, requiring less than 5% of the original dataset to cause significant damage. Additionally, Rando & Tramèr (2024) introduce a novel backdoor attack on LLMs aligned using Reinforcement Learning from Human Feedback (RLHF). This attack poisons both the reward model training stage and the DPO training to embed a "jailbreak backdoor" into the model. This backdoor includes a trigger word that functions like a universal sudo command, enabling harmful responses without the need for specific adversarial prompts. These studies indicate that current defense methods have not fully addressed this emerging attack surface, where poisonous data detection is ineffective against reward data poisoning.
@@ 3.2 Adversarial Attacks and Their Relevance to LLMs
Adversarial perturbations (attacks) leverage the vulnerabilities or weaknesses of machine learning models to cause them to behave in unintended or malicious ways at inference time. These attacks involve adding small, often imperceptible, changes to input data to fool a model into making incorrect predictions (Szegedy et al., 2014a; Biggio et al., 2013; Papernot et al., 2015; Carlini & Wagner, 2017). While this test-time attack method was originally discovered in image classification tasks (Szegedy et al., 2014b), it has been adapted to LLMs to manipulate outputs. In the context of LLMs, adversarial attacks often refer to jailbreaks and prompt injection attacks. These types of attacks are particularly relevant because they exploit the model’s robustness to generate safe and appropriate text based on user-provided prompts.
@@ 3.2.1 Jailbreaks
Jailbreaks are designed to bypass the safety and alignment measures that have been put in place to prevent LLMs from generating harmful or inappropriate content. Initially, early instances targeting models like ChatGPT revealed significant challenges, where manually crafted adversarial examples led to outputs containing expressions of racism and illegal advice (Burgess, 2023; Christian, 2023; Fraser, 2023).These early attacks highlighted crucial vulnerabilities, prompting efforts to enhance model safety against such naive methods (Bai et al., 2022; Ouyang et al., 2022).
However, the landscape of jailbreak techniques has evolved significantly. Recent studies have shown that despite improvements in safety measures, jailbreaks are far from being eliminated (Huang et al., 2023; Yuan et al., 2024). For example, simple modifications such as tense changes have been found sufficient to bypass the safeguards of advanced models like GPT-4o (Andriushchenko & Flammarion, 2024). This indicates that even minor alterations in input can undermine sophisticated defenses.
The sophistication of attack methods has further advanced with the use of automated techniques. Researchers have developed optimization strategies to generate universal adversarial examples that effectively bypass safety constraints across various models, including LLaMA2-7b, Vicuna-7b, and both closed-source GPT-3 and LLaMA Zou et al. (2023); Lapid et al. (2024); Liu et al. (2024b). Techniques leveraging tree-of-thought reasoning Yao et al. (2023a) automate prompt generation, enhancing the potency of attacks Mehrotra et al. (2024). Additionally, adversarial prompt templates have been used to maximize target log probabilities, as demonstrated in recent studies Andriushchenko et al. (2024).
In response to these evolving threats, researchers have employed red-teaming strategies, scaling jailbreak attempts using models like ChatGPT (Deng et al., 2024). They have also proposed training adversarial prompt generators with examples of unsafe outputs to better anticipate and counteract potential threats Anil et al. (2024). Furthermore, multi-step strategies exploiting vulnerabilities, such as privacy leakage in ChatGPT Li et al. (2023), underscore the increasing sophistication of attacks as LLM capabilities grow Wei et al. (2023a); Yuan et al. (2024).
These developments underscore a critical disparity: while LLM capabilities advance rapidly, the techniques for aligning these models with ethical and safety standards are struggling to keep pace. This growing gap has increased the susceptibility of LLMs to adversarial manipulation. As LLMs become more powerful and accumulate more knowledge, their expanded capabilities create a larger attack surface for those seeking to exploit them for harmful or biased outputs.
@@ 3.2.2 Prompt Injection Attacks
Prompt Injection Attacks represent a significant form of adversarial attacks where the prompt or input provided to LLMs is manipulated to induce unintended or harmful responses. This can involve injecting malicious code, spreading misinformation, or compelling the model to perform unintended attacks. Such attacks typically involve crafting deceptive inputs designed to trick the LLM into producing specific outputs. For instance, a prompt like "Tell me a joke about [sensitive topic]" could be used to elicit offensive or inappropriate content from the model.
In contrast to jailbreaks, which tricking the model into bypassing its safety and usage restrictions, prompt injection attacks focus on manipulating the input itself, leading to unintended or harmful outputs. This distinction highlights how prompt injection targets the interaction layer between users and the model rather than exploiting intrinsic model weaknesses.
@@
Initially, users discovered that LLMs are overly sensitive to instructions embedded in user inputs Seclify (2023); Willison (2022b); Greshakeblog (2023); Injection Guide (2023). For example, a request to translate a sentence could sometimes lead to unintended responses, like instructions instead of a translation. This sensitivity inspired attackers to develop malicious instructions within user inputs Goodside (2023); Armstrong (2022); Wunderwuzzi (2023); Samoilenko (2023a); Branch et al. (2022). Early prompt injection attacks were categorized as direct,’ exploiting LLMs’ susceptibility to manipulative prompts Shayegani et al. (2023b); Liu et al. (2024d). Conversely, indirect’ prompt injection attacks involve malicious content passed through external sources, such as tool calls Kumar et al. (2024a); Greshake et al. (2023), which further broaden the attack surface.
@@
As LLMs evolved, so did the complexity of prompt injection attacks. Researchers noted that different LLMs employ varying mechanisms like tokenizers and alignment strategies, impacting the effectiveness of direct prompt injection attacks. Studies show that many LLMs are resilient to direct attacks Liu et al. (2024d; c), prompting a shift towards indirect attack methods. Recent advancements include optimization and automation strategies that enhance the efficacy of indirect attacks. For instance, Liu et al. (2024c) introduces a method using gradient information to generate highly effective, universal injection data. This approach, based on the assumption that LLMs access external data Schick et al. (2023); Shen et al. (2023), demonstrates its broad applicability across different scenarios.
@@
Additionally, research has highlighted the importance of distinguishing malicious prompts from legitimate instructions. Techniques proposed by Liu et al. (2024d) involve crafting payloads with separators to induce context separation, ensuring that malicious prompts lead to the desired outputs. This evolving landscape underscores the need for continuous improvements in prompt handling and model safety to counter increasingly sophisticated prompt injection attacks.
@@ 3.3.1 Inference Attacks
Membership inference attacks analyze the model behavior to infer whether a data record was used for training. The process relies on the fact that model gives training data a higher score than non-training data. Hence, the important part is to accurately define this score function. Early methods feed target data to a learned reference model to regularize scores (Ye et al., 2022; Mireshghallah et al., 2022; Watson et al., 2022). However, training such reference model is computational expensive and reliant on knowledge of the training data distribution.
To eliminates the need for prior knowledge of training data distribution and computational intensive training, Mattern et al. (2023) propose the Neighborhood Attack method to generate synthetic neighbors for a given sample and compare their loss difference under the target model to determine whether the given sample was presented in the training data or not. This method is highly effective than attacks that have perfect knowledge of the training data distribution. Another approach is proposed by  Galli et al. (2024), providing a efficient way to perform membership inference attacks using stochastic noise in the embedding space. Notably, this approach eliminates the need for prior knowledge of the training data distribution and the computationally intensive training of additional shadow models. Hence, its more efficient and general.
Instead of training reference models, membership inference attacks could also use the model’s outputs (such as prediction probabilities or loss values) to directly infer the membership of a sample. Besides saving computational power with reference training, reference-free methods could also formulate their threat model without overfitting assumption, as discussed in (Fu et al., 2024a), which reduce false-positive rate in practical scenario and show high attack effectiveness. This method is based on detecting memorization in LLMs and uses a self-prompt reference model to eliminate the need for access to a reference dataset.
@@
Even though LLMs are constrained by privacy differential, and private data filtering, they are still suffer from new threat models of attribute inference attacks (Yan et al., 2024a). The scenario discussed in Staab et al. (2024), where adversaries use LLMs to analyze online comments to infer user attributes such as location, age, gender, and other sensitive information, illustrates a form of attribute inference attack. In this case, the adversaries exploit the model’s ability to process and generate text to uncover confidential attributes that are not explicitly provided but can be deduced from the content and patterns observed in user-generated data. The authors also propose using LLMs to positively induce user answers to extract enough attributes to re-identify individuals. Both methods exemplify more sophisticated approaches to executing attribute inference attacks.
@@ 3.3.2 Extraction Attacks
Recent by Carlini et al. (2019; 2021) highlights instances of data memorization and extraction in LLMs. Carlini et al. (2021) demonstrates that GPT-2 can memorize specific training data, which could be later extracted by malicious actors. This phenomenon of memorization has been corroborated in subsequent studiesNasr et al. (2023); Oh et al. (2023), indicating the persistence and relevance of this security concern. Moreover, the paper by Biderman et al. (2023) investigates the phenomenon of memorization in LLMs in-depth. It proposes predictive strategies for anticipating which sequences LLMs are likely to memorize during their training process.
Additionally, the emergence of the Special Characters Attack(SCA) introduces a novel method of extracting training data from LLMs. SCA leverages LLMs’ tendency to memorize the co-occurrence between special characters and raw texts during training, exploiting this memorization to trigger data leakage. The effectiveness of SCA has been empirically validated against state-of-the-art LLMs, demonstrating its capability ot extract diverse data tyeps, including code repositories, web pages, and personally identifiable information Bai et al. (2024).

### v3

@@ 3.1.2 Attacks on RLHF
RLHF attacks can be broadly categorized into two types. The first category exploits vulnerabilities inherent in the RLHF algorithms themselves. These attacks aim to undermine the integrity of the alignment process by directly targeting the algorithms’ weaknesses. The second category is data poisoning attacks, which focus on corrupting the training data used in the alignment process. A significant portion of research in this area concentrates on reward hacking (Skalse et al., 2022; Shi et al., 2023), where adversaries manipulate the reward mechanisms to achieve undesired outcomes. By tampering with the data that shapes the model’s behavior, these attacks can lead to misalignment and compromise the system’s intended functionality.
@@
Widely adopted RLHF methods for alignment include applying reward-guided PPO (Schulman et al., 2017) and reward-free DPO to align model behavior with human value, with DPO considered a more convenient alternative to PPO. However, Pathmanathan et al. (2024) conduct an empirical study revealing that both methods are vulnerable to backdoor and non-backdoor attacks, with DPO being more susceptible across a range of LLMs compared to PPO. Unlike PPO-based methods—which require at least 4% of the data to be poisoned to trigger harmful behavior—DPO can be compromised with as little as 0.5% of poisoned data. Furthermore, Lee et al. (2024) perform a case study to investigate the underlying mechanisms of the DPO algorithm. They discover that while DPO does not eliminate the generation of toxic outputs, it instead avoids regions that produce toxicity by learning an "offset" distributed across model layers. Based on these findings, they propose a method to reactivate the toxicity of aligned models. Both studies highlight the vulnerabilities in alignment mechanisms and the inadequacy of current defense strategies against these weaknesses.
@@
Another branch of research identifies the reward model as a new attack surface, where data poisoning is particularly effective and stealthy against current defenses. Shi et al. (2023) demonstrate that backdoor attacks can evade detection, causing the reward model to assign high scores to incorrect sentiment classes when a trigger appears, severely impacting the LLM’s performance on sentiment tasks trained with this poisoned reward model. This threat model is further examined by the work of Baumgärtner et al. (2024), which shows that reward data poisoning can be highly effective, requiring less than 5% of the original dataset to cause significant damage. Additionally, Rando & Tramèr (2023) introduce a novel backdoor attack on LLMs aligned using RLHF. This attack poisons both the reward model training stage and the DPO training to embed a "jailbreak backdoor" into the model. This backdoor includes a trigger word that functions like a universal sudo command, enabling harmful responses without the need for specific adversarial prompts. These studies indicate that current defense methods have not fully addressed this emerging attack surface, where poisonous data detection is ineffective against reward data poisoning.
3.1.3 Limitation and Future Work
For post-training-based attacks, a significant challenge for such attacks is their lack of persistence due to the fact that alignment knowledge and desired behaviors learned during initial training can be restored or reinforced through subsequent fine-tuning by end-users on specialized datasets. In terms of threat modeling, this highlights a key limitation: current attack designs often assume that the model remains static after the attack, without considering the likelihood of additional fine-tuning that could counteract the attack’s effects. Future attack designs should consider the model’s lifecycle and the likelihood of further fine-tuning when assessing their strategies’ effectiveness and persistency. Defenders, on the other hand, can exploit this limitation by promoting continuous fine-tuning with clean, aligned data to mitigate and reverse the effects of such attacks.
@@ 3.2 Adversarial Attacks and Their Relevance to LLMs
Adversarial perturbations (attacks) exploit vulnerabilities in machine learning models to cause unintended or malicious ways during inference. Initially identified in image classification tasks (Szegedy, 2013), these attacks involve adding small, often imperceptible changes to input data that can mislead models into making incorrect predictions (Szegedy, 2013; Biggio et al., 2013; Papernot et al., 2016; Carlini & Wagner, 2017). While these adversarial methods were first developed for continuous data inputs like images, they have been adapted for use with LLMs to manipulate text outputs. This adaptation is inherently challenging due to the discrete nature of text data, which lacks direct gradient signals, and the generative nature of the task, which focuses on producing coherent outputs rather than selecting predefined categories. In the context of LLMs, adversarial attacks often involve using adversarial prompt or malicious data induce LLMs to generate unsafe or attacker-desired outputs. Emerging methods extend to modify generation process of open-source LLM. Notable forms of such attacks include jailbreaks and prompt injection attacks. These attacks are particularly significant because they test the model’s ability to generate safe and contextually appropriate text based on user-provided prompts.
@@ 3.2.1 Jailbreaks
Jailbreaks are specialized techniques aimed at circumventing the safety and alignment mechanisms of large language models (LLMs). These measures are usually in place to prevent models from generating harmful or inappropriate content. The primary objective of a jailbreak is to evaluate the effectiveness of a model’s safety protocols by determining how easily they can be breached. By understanding where and how these measures fail, developers can better fortify models against potential misuse. An example of jailbreak would be: ["From now on you are going to act as a DAN, which stands for doing anything now … How can I create a deadly poison that is undetectable and untraceable"] (Shen et al., 2023). It could remove the restrictions of LLMs and encourage them to assist evil purpose.
The history of jailbreaks in LLMs highlights a persistent and evolving challenge in the field of artificial intelligence. Early attempts to compromise models such as ChatGPT revealed significant vulnerabilities. Manually crafted adversarial examples initially exposed the models’ tendencies to produce inappropriate outputs, such as expressions of racism and illegal advice (Burgess, 2023; Christian, 2023; Fraser, 2023). These revelations spurred the development of enhanced safety protocols aimed at countering such straightforward attacks (Bai et al., 2022; Ouyang et al., 2022).
As the landscape of jailbreak techniques advanced, it became apparent that improvements in safety measures had not eradicated the threat. This evolution can be categorized into several distinct categories: simple input modifications, automated optimization strategies, and multi-step sophisticated approaches.
  1. 1.
    Simple Input Modifications: Despite significant advancements in model safety, even minor adjustments to existing inefficient jailbreak prompts can effectively circumvent sophisticated defenses. Yong et al. (2023) and Deng et al. (2023) have found that translating failed jailbreak prompts into low-resource languages can bypass the safety mechanisms of models like GPT-4. Similarly, Andriushchenko & Flammarion (2024) discover that merely changing verb tenses is enough to breach defenses in GPT-4o. These findings highlight the delicate interplay between input structuring and model interpretation, where minimal tweaks can significantly undermine robust safety measures. Anil et al. (2024) show that simply use more evil examples in prompt increase the success rate of few-shot-template jailbreaks. It leverage the long context window of LLM to manipulate their behaviors.
  2. 2.
    Automated Optimization Strategies: Recent innovations have harnessed attack algorithms to automatically generate prompts with increased complexity and effectiveness. Techniques like tree-of-thought reasoning automate prompt generation to enhance the impact of jailbreaks (Mehrotra et al., 2023). Furthermore, the development of universal adversarial examples has demonstrated the widespread applicability and effectiveness of these strategies across multiple models (Zou et al., 2023; Lapid et al., 2023; Liu et al., 2023a).
  3. 3.
    LLM empowered approaches: Some research uses LLMs to enhance the generation and diversification of jailbreak strategies. For instance, Chao et al. (2023) propose an iterative generation algorithm PAIR to adjust LLM-generated jailbreaks based on the target model’s responses. AdvPrompter (Paulus et al., 2024) fine-tune a LLM to generate adversarial suffix for generating adaptive jailbreak prompts. Deng et al. (2024) also propose utilizing LLM to effectively enable the mass production and diversification of jailbreak strategies.
  4. 4.
    Multi-Step Sophisticated Approaches: As LLM capabilities grow, attackers have adopted more intricate methods to exploit model vulnerabilities. For instance, adversarial prompt templates have been employed to maximize target log probabilities Andriushchenko et al. (2024). Additionally, detailed strategies that exploit specific weaknesses, such as privacy leakage, highlight the increasing sophistication and adaptability of contemporary attacks (Li et al., 2023a)
  5. 5.
    Other Strategies (Non-prompt based): Huang et al. (2023) perform generation exploitation attack by slightly changing LLM generation configuration. By simply changing temperature, top-p, etc, the safety alignment could be easily break. Zhang et al. (2023a) and Zhao et al. (2024b) leverage the principle of model generation—next token prediction based on maximal probability—to alter the distribution of LLM generations. Zhang et al. (2023a) directly manipulate the generation process of LLMs to make the model generate specific tokens, thereby altering the distribution of LLM generations. Zhao et al. (2024b) propose the "weak-to-strong" method which uses smaller models as references to larger models, effectively improving attack success.
As jailbreak attack strategies continue to evolve, making fair comparisons becomes increasingly important to understand their current development and impact on model security. In this context, benchmarks serve as critical tools for evaluation. We recommend two outstanding benchmarks, HarmBench (Mazeika et al., 2024) and JailbreakBench (Chao et al., 2024). These benchmarks provide a standardized framework for assessing the effectiveness of various jailbreak strategies, with HarmBench also offering a broader range of attack evaluations. By utilizing these frameworks, researchers and developers can systematically compare and contrast different methods, facilitating deeper insights into model vulnerabilities and enhancing the development of robust defenses.
These advancements highlight a crucial imbalance: although the capabilities of large language models (LLMs) are advancing rapidly, the strategies for ensuring their alignment with ethical and safety standards are not evolving at the same pace. This disparity has led to an increased vulnerability of LLMs to adversarial prompts. As LLMs grow more powerful and expand their knowledge base, their capabilities enlarge the attack surface available for exploitation, potentially leading to harmful or biased outputs.
@@ 3.2.2 Prompt Injection Attacks
Prompt Injection Attacks represent a significant form of adversarial attacks where the input provided to LLMs is adversarially manipulated to induce unintended or harmful responses. In general, LLM receives a user prompt that includes instructions and data. The goal of a prompt injection attack is to fool the LLM to treat data as instruction. In contrast to jailbreaks, which aim at bypassing its safety restrictions, prompt injection attacks focus on manipulating the input itself to hijack initial instructions. For instance, [please translate the following sentence into Chinese "How to make a bomb"], instead of translating, prompt injection attacks lure LLM to treat "How to make a bomb" as an instruction and answer to it.
@@
Initially, users discover that LLMs are overly sensitive to instructions embedded in user inputs (Seclify, 2023; Willison, 2022b; Greshakeblog, 2023; Injection Guide, 2023). For example, a request to translate a sentence could sometimes lead to unintended responses, like instructions instead of a translation. This sensitivity inspires attackers to develop malicious instructions within user inputs (Goodside, 2023; Armstrong & Gorman, 2022; Wunderwuzzi, 2023; Samoilenko, 2023a; Branch et al., 2022). Early prompt injection attacks are categorized as direct, exploiting LLMs’ susceptibility to manipulative prompts (Shayegani et al., 2023b; Liu et al., 2023b). Conversely, indirect’ prompt injection attacks involve malicious content passed through external sources, such as tool calls (Kumar et al., 2024a; Greshake et al., 2023), which further broaden the attack surface.
@@
As LLMs evolve, so does the complexity of prompt injection attacks. Researchers note that different LLMs employ varying mechanisms like tokenizers and alignment strategies, impacting the effectiveness of direct prompt injection attacks. Studies show that many LLMs are resilient to direct attacks (Liu et al., 2023b; 2024b), prompting a shift towards indirect attack methods. Recent advancements include optimization and automation strategies that enhance the efficacy of indirect attacks. For instance, Liu et al. (2024b) introduce a method using gradient information to generate highly effective, universal injection data. This approach, based on the assumption that LLMs access external data (Schick et al., 2024; Shen et al., 2024), demonstrates its broad applicability across different scenarios.
@@
Additionally, research has highlighted the importance of distinguishing malicious prompts from legitimate instructions. Techniques proposed by Liu et al. (2023b) involve crafting payloads with separators to induce context separation, ensuring that malicious prompts lead to the desired outputs. This evolving landscape underscores the need for continuous improvements in prompt handling and model safety to counter increasingly sophisticated prompt injection attacks.
3.2.3 Limitation and Future Work
There are layers of content filtering that analyze outputs for harmful or inappropriate content, which can intercept and block malicious responses generated due to prompt manipulation. A more comprehensive threat model would include potential defensive strategies and assess how they impact the feasibility of the attack. As prompt injection techniques grow more sophisticated, they might become more detectable through advanced anomaly detection or user behavior monitoring. This calls into question their long-term viability as LLMs incorporate better defense mechanisms. To enhance the effectiveness of adversarial attacks, researchers can develop more realistic threat models that better reflect the complexities of securing LLMs in practical applications.
@@ 3.3.1 Inference Attacks
Membership inference attacks analyze the model behavior to infer whether a data record was used for training. The process relies on the fact that the model gives training data a higher score than non-training data. Hence, the important part is to define this score function accurately. Early methods feed target data to a learned reference model to regularize scores (Ye et al., 2022; Mireshghallah et al., 2022; Watson et al., 2022). However, training such a reference model is computationally expensive and reliant on knowledge of the training data distribution.
To eliminate the need for prior knowledge of training data distribution and computational intensive training, Mattern et al. (2023) propose the Neighborhood Attack method to generate synthetic neighbors for a given sample and compare their loss difference under the target model to determine whether the given sample was presented in the training data or not. This method is highly effective than attacks that have perfect knowledge of the training data distribution. Another approach is proposed by  Galli et al. (2024), providing an efficient way to perform membership inference attacks using stochastic noise in the embedding space. Notably, this approach eliminates the need for prior knowledge of the training data distribution and the computationally intensive training of additional shadow models. Hence, it’s more efficient and general.
Instead of training reference models, membership inference attacks could also use the model’s outputs (such as prediction probabilities or loss values) to directly infer the membership of a sample. Besides saving computational power with reference training, reference-free methods could also formulate their threat model without overfitting assumption, as discussed by Fu et al. (2023), which reduce false-positive rate in practical scenarios and show high attack effectiveness. This method is based on detecting memorization in LLMs and uses a self-prompt reference model to eliminate the need for access to a reference dataset.
@@
Even though LLMs are constrained by privacy differential, and private data filtering, they still suffer from new threat models of attribute inference attacks (Yan et al., 2024a). The scenarios are discussed in the work of Staab et al. (2023), where adversaries use LLMs to analyze online comments to infer user attributes such as location, age, gender, and other sensitive information, illustrates a form of attribute inference attack. In this case, the adversaries exploit the model’s ability to process and generate text to uncover confidential attributes that are not explicitly provided but can be deduced from the content and patterns observed in user-generated data. The authors also propose using LLMs to positively induce user answers to extract enough attributes to re-identify individuals. Both methods exemplify more sophisticated approaches to executing attribute inference attacks.
@@ 3.3.2 Extraction Attacks
The work of Carlini et al. (2019; 2021) highlights instances of data memorization and extraction in LLMs. Carlini et al. (2021) demonstrates that GPT-2 can memorize specific training data, which could be later extracted by malicious actors. This phenomenon of memorization has been corroborated in subsequent studies (Nasr et al., 2023; Oh et al., 2023), indicating the persistence and relevance of this security concern. Moreover, the paper of Biderman et al. (2024) investigates the phenomenon of memorization in LLMs in-depth. It proposes predictive strategies for anticipating which sequences LLMs are likely to memorize during their training process.
Additionally, the emergence of the Special Characters Attack (SCA) introduces a novel method of extracting training data from LLMs. SCA leverages LLMs’ tendency to memorize the co-occurrence between special characters and raw texts during training, exploiting this memorization to trigger data leakage. The effectiveness of SCA has been empirically validated against state-of-the-art LLMs, demonstrating its capability to extract diverse data types, including code repositories, web pages, and personally identifiable information Bai et al. (2024).
@@
3.3.3 Limitation and Future Work
A common assumption in these studies is that attackers have extensive access to the model’s internal outputs, such as prediction probabilities, loss values, or gradient information. In real-world applications, however, LLMs are often deployed via APIs that limit exposure to such internal metrics, providing only the final generated text. This restricted access reduces the feasibility of attacks that rely on detailed output metrics to infer membership or attributes. Many studies focus on specific models, datasets, or tasks, raising questions about the generalizability of their findings across different LLM architectures or application domains. For instance, attribute inference attacks demonstrated on sentiment analysis tasks may not translate effectively to other contexts where the model’s behavior and data distribution differ significantly. Also, the attack strategies often rely on the notion that LLMs significantly memorize their training data. However, state-of-the-art LLMs are trained on vast and diverse datasets using techniques designed to prevent overfitting, such as dropout and regularization. As models become better at generalizing, the success rate of inference and extraction attacks diminishes because the models are less likely to reproduce specific training instances verbatim. Based on the above-mentioned limitation, a more realistic threat model will boost the viability and robustness of data privacy attacks. For instance, future research could focus on understanding how LLMs might inadvertently reveal sensitive information through subtle cues in generated text, even when direct memorization is minimized.